# Amoeba Monte Carlo algorithms for random trees with controlled branching activity: efficient trial move generation and universal dynamics

### v1

@@ I Introduction
In the present paper, we explore how to efficiently treat the case where the average number of linear segments between branch-points is controlled by a branching activity or a correponding chemical potential $\mu$ [62, 60, 63] favoring or suppressing the presence of branch-points. On first sight, one might expect this to be of little consequence for the efficiency of the Amoeba algorithm. After all, there are even more end nodes to choose from in the limit of strong branching, while one might hope to recover the equally good performance of the reptation algorithm when dealing with essentially linear chains in the opposite limit of no branching. However, closer inspection reveals a strong decrease of the acceptance probability for the classical Amoeba moves. For example, when branching is suppressed, most attempts to attach the severed segment to one of the inner nodes of the truncated linear chain are rejected. It turns out that the situation is even worse for strongly branching trees with an even number of nodes, which need to pass through an exponentially suppressed excited state containing a linear chain segment. Here, we show how to modify the Amoeba algorithm to render its efficiency independent of $\mu$ and present a detailed analysis of the relaxation dynamics.
@@ II.3 Daoud-Joanny theory for ideal random trees
$\langle n_{3}(N_{b})\rangle=\frac{N_{b}}{2}\frac{\partial{\log(\mathcal{Z}}_{N_{b}})}{\partial N_{b}}\\ \simeq\left\{\begin{array}[]{lc}\frac{(\lambda N_{b})^{2}}{2}\,,&\lambda N_{b}\ll 1\\ \\ \lambda N_{b}\,,&\lambda N_{b}\gg 1\end{array}\right.\,.$ (10)
@@ III Monte Carlo methods
    The double-leaf Amoeba Amoeba algorithm (Sec. III.3) employs additional trial moves affecting peripheral branch-points connecting two leaves to the rest of the tree. After removing both leaves from a randomly selected branch-point of this type (converting it into a leaf), they are attached to a randomly chosen leaf (converting it into a peripheral branch-point).
@@ III.1 Original Amoeba algorithm
${\rm acc}_{|i\rangle\rightarrow|f\rangle}=\min\left\{1,\,\,\,\frac{n_{1}(|f\rangle)}{n_{1}(|i\rangle)}\,e^{\beta\mu(n_{3}(|f\rangle)-n_{3}|i\rangle)}\right\}\,.$ (12)
@@ III.2 Semi-kinetic Amoeba algorithm
${\rm acc}_{|i\rangle\rightarrow|f\rangle}=\min\left\{1,\,\frac{w(|f\rangle)\,\alpha(|f\rangle\rightarrow|t\rangle)\,\alpha(|t\rangle\rightarrow|i\rangle)}{w(|i\rangle)\,\alpha(|i\rangle\rightarrow|t\rangle)\,\alpha(|t\rangle)\rightarrow|f\rangle)}\right\}\,.$ (13)
@@
$\alpha(|i\rangle\rightarrow|t\rangle)=\frac{1}{n_{1}(|i\rangle)}\,,\,\,\,\alpha(|f\rangle)\rightarrow|t\rangle)=\frac{1}{n_{1}(|f\rangle)}\,.$ (14)
@@ V.3 A universal tree regime in the Amoeba dynamics
Importantly, the color coding reveals that the deviations from ordinary diffusion are entirely controlled by the number
$n_{\rm lin}=2n_{3}+1$ (26)
of linear segments in the trees. Linear chains with $n_{\rm lin}=1$ exhibit the expected time-independent center-of-mass diffusion. For branched trees, the larger $n_{\rm lin}\gg 1$, the longer the anomalous diffusion regime and the larger the slowdown of the asymptotic diffusion.
@@
where $\tau_{0}$ is a microscopic time from the point of view of the tree dynamics. In Eq. (27) the intermediate time scale $\tau_{1}$ separates (i) the early regime during which the algorithm predominantly performs mass exchange between the outer linear segments of the trees from (ii) the tree regime controlled by the destruction and creation of branch-points, during which the tree centers-of-mass exhibit anomalous diffusion. Notice that the two proposed scaling behaviors for $\tau_{1}$ and $\tau_{\rm eq}$ in Eqs. (28) and (29) follow from the mentioned observation that $n_{\rm lin}$ controls the deviations of $g_{3}$ from pure diffusion. In particular, the two characteristic time scales converge to the standard reptation expression in the absence of branch points: $\lim_{n_{\rm lin}\rightarrow 1}\tau_{1}=\lim_{n_{\rm lin}\rightarrow 1}\tau_{\rm eq}=N^{2}\tau_{0}$.
@@ VI Conclusion and outlook
Acknowledgements – PHWvdH acknowledges the kind hospitality of the ENS-Lyon.

### v2

@@ I Introduction
In the present paper, we explore how to efficiently treat the case where the average number of linear segments between branch-points is controlled by a branching activity or a corresponding chemical potential $\mu$ [62, 60, 63] favoring or suppressing the presence of branch-points. On first sight, one might expect this to be of little consequence for the efficiency of the Amoeba algorithm. After all, there are even more end nodes to choose from in the limit of strong branching, while one might hope to recover the equally good performance of the reptation algorithm when dealing with essentially linear chains in the opposite limit of no branching. However, closer inspection reveals a strong decrease of the acceptance probability for the classical Amoeba moves. For example, when branching is suppressed, most attempts to attach the severed segment to one of the inner nodes of the truncated linear chain are rejected. It turns out that the situation is even worse for strongly branching trees with an even number of nodes, which need to pass through an exponentially suppressed excited state containing a linear chain segment. Here, we show how to modify the Amoeba algorithm to render its efficiency independent of $\mu$ and present a detailed analysis of the relaxation dynamics.
@@ II.3 Daoud-Joanny theory for ideal random trees
$\langle n_{3}(N_{b})\rangle=\frac{N_{b}}{2}\frac{\partial{\log(\mathcal{Z}}_{N_{b}})}{\partial N_{b}}\simeq\left\{\begin{array}[]{lc}\frac{(\lambda N_{b})^{2}}{2}\,,&\lambda N_{b}\ll 1\\ \\ \lambda N_{b}\,,&\lambda N_{b}\gg 1\end{array}\right.\,.$ (10)
@@ III Monte Carlo methods
    The double-leaf Amoeba algorithm (Sec. III.3) employs additional trial moves affecting peripheral branch-points connecting two leaves to the rest of the tree. After removing both leaves from a randomly selected branch-point of this type (converting it into a leaf), they are attached to a randomly chosen leaf (converting it into a peripheral branch-point).
@@ III.1 Original Amoeba algorithm
${\rm acc}_{|i\rangle\rightarrow|f\rangle}=\min\left\{1,\,\,\,\frac{n_{1}(|f\rangle)}{n_{1}(|i\rangle)}\,e^{\beta\mu(n_{3}(|f\rangle)-n_{3}(|i\rangle))}\right\}\,.$ (12)
@@ III.2 Semi-kinetic Amoeba algorithm
${\rm acc}_{|i\rangle\rightarrow|f\rangle}=\min\left\{1,\,\frac{w(|f\rangle)\,\alpha(|f\rangle\rightarrow|t\rangle)\,\alpha(|t\rangle\rightarrow|i\rangle)}{w(|i\rangle)\,\alpha(|i\rangle\rightarrow|t\rangle)\,\alpha(|t\rangle\rightarrow|f\rangle)}\right\}\,.$ (13)
@@
$\alpha(|i\rangle\rightarrow|t\rangle)=\frac{1}{n_{1}(|i\rangle)}\,,\,\,\,\alpha(|f\rangle\rightarrow|t\rangle)=\frac{1}{n_{1}(|f\rangle)}\,.$ (14)
@@ V.3 A universal tree regime in the Amoeba dynamics
Importantly, the color coding reveals that the deviations from ordinary diffusion are entirely controlled by the mean number
$\langle n_{\rm lin}\rangle=2\langle n_{3}\rangle+1$ (26)
of linear segments in the trees. Linear chains with $\langle n_{\rm lin}\rangle=1$ exhibit the expected time-independent center-of-mass diffusion. For branched trees, the larger $\langle n_{\rm lin}\rangle\gg 1$, the longer the anomalous diffusion regime and the larger the slowdown of the asymptotic diffusion.
@@
where $\tau_{0}$ is a microscopic time from the point of view of the tree dynamics. In Eq. (27) the intermediate time scale $\tau_{1}$ separates (i) the early regime during which the algorithm predominantly performs mass exchange between the outer linear segments of the trees from (ii) the tree regime controlled by the destruction and creation of branch-points, during which the tree centers-of-mass exhibit anomalous diffusion. Notice that the two proposed scaling behaviors for $\tau_{1}$ and $\tau_{\rm eq}$ in Eqs. (28) and (29) follow from the mentioned observation that $\langle n_{\rm lin}\rangle$ controls the deviations of $g_{3}$ from pure diffusion. In particular, the two characteristic time scales converge to the standard reptation expression in the absence of branch points: $\lim_{\langle n_{\rm lin}\rangle\rightarrow 1}\tau_{1}=\lim_{\langle n_{\rm lin}\rangle\rightarrow 1}\tau_{\rm eq}=N^{2}\tau_{0}$.
@@ VI Conclusion and outlook
Acknowledgements – PHWvdH acknowledges financial support from PNRR_M4C2I4.1._DM351 funded by NextGenerationEU and the kind hospitality of the ENS-Lyon. AR acknowledges financial support from PNRR Grant CN_00000013_CN-HPC, M4C2I1.4, spoke 7, funded by Next Generation EU.